# Machine Unlearning Through Fine-Grained Model Parameters Perturbation

### v3

@@ IV-C2 Determination of $K$ value for Top-K
We aim to achieve the effect of inexact machine unlearning by minimal number of parameter perturbed, to investigate the optimal value of $K$ for Top-K unlearning strategy, we conduct experiments with different values of $K$ within the range of 0 to 50 when the unlearning data ratio varies from 5% to 20%. Following the above settings, we apply Top-K to four models over 100 epochs. Since machine unlearning aims to eliminate the influence of unlearning data while trying to not hurt the performance of remaining data, we calculate the accuracy difference between $Acc_{RE}$ and $Acc_{UL}$. As illustrated in Fig.5, the $K$ value corresponding to the maximum accuracy difference is identified as the optimal. As shown in Tab.I, considering the optimal $K$ values under four unlearning data ratios, we set $K=45$ for ResNet18 and DenseNet, $K=30$ for VGG, and $K=50$ for GoogLeNet.
As shown in Fig.5, although we have chosen the $K$ corresponding to the maximum accuracy difference, in actual experiments, the selection of $K$ values fluctuating within the range of 0 to 50 does not bring significant variability to the Top-K unlearning. For example, in Fig.5(a), for ResNet18, setting K=25 versus 45 does not make a substantial difference. Additionally, measuring the difference in accuracy is a sufficient but not necessary condition for the effectiveness of machine unlearning. Therefore, additional efforts or computational costs to determine a very precise $K$ value are unnecessary; we can roughly set it to a value such as 45.
IV-C3 Effectiveness of Top-K and Random-k
Fig.1 further shows the accuracy after applying the Random-k on the CIFAR-10 for 50 epochs. The accuracy of $D_{RE}$ is significantly higher than that of $D_{UL}$. It can be observed that the application of the Top-K achieves a 95% $Acc_{RE}$ across the four models, indicating a strong memory retention capability. Meanwhile, the accuracy difference between unlearning model on the $D_{RE}$ and the $D_{UL}$ (both in Fig.6 and Fig.1) shows that both the Top-K and Random-k achieve the unlearning effect more or less.
@@ IV-C3 Effectiveness of Top-K and Random-k
$D_{UL}/D$ Strategy $Acc_{UL}$ $Acc_{RE}$ FR MRR Similarity Unlearn Time
5% Top-K 80.01% 97.61% 19.97% 97.63% 88.37% 76.97s
@@
Tab.II provides a more comprehensive overview of the effects of different unlearning strategies at various unlearning ratios. Across different unlearning ratios, Top-K consistently exhibits the lowest $Acc_{UL}$ values (80.14%, 75.48%, and 75.46%), along with corresponding forgetting rates (19.86%, 24.52%, and 24.54%). This suggests that, relative to other strategies, Top-K achieves the most effective reduction of the impact of $D_{UL}$ on the model when the unlearning ratio belongs to 10%, 15%, 20%. However, for the 5% unlearning ratio under Random-k, its corresponding $Acc_{RE}$ and MRR of 90.07% noticeably lag behind the results of other strategies. We believe this disparity is due to the instability of the strategy that employs random parameter perturbation, leading to inconsistent unlearning effects. In the meanwhile, across both $Acc_{RE}$ and MRR, the performance of $M_{UL}$ from nearly all unlearning strategies is comparable to that of retraining models, achieving close to 100% accuracy. As for similarity, there isn’t a significant distinction among different strategies, but this does indicate that model indistinguishability is maintained across diverse strategies.
@@
To further illustrate the acceleration effect of unlearning strategies on different models compared to the retraining method, we also conduct machine unlearning tasks on VGG, DenseNet, and GoogLeNet. Fig.7 shows the acceleration ratio of each unlearning strategy compared to retraining under different models and unlearning ratios. The acceleration effect of Top-K and Random-k is better than other strategies. Taking Fig.7(a) as an example, when the unlearning ratio is 5%, Top-K can achieve an acceleration of more than 15x, while Random-k achieves an acceleration close to 9x. The acceleration from EU-K and CF-K methods is not significant. As the unlearning ratio increases, the acceleration of Top-K and Random-k slows down. This is because, as more data is to be forgotten, the degree of modification to the model deepens. In conclusion, we can ascertain that the Top-K achieves optimal unlearning performance with minimal parameter perturbed, maintains considerable memory retention capability, and achieves the fastest acceleration effect.
@@ IV-C4 Mixed Top-K and Random-k
In Random-k, parameters are chosen randomly to be perturbed, while in Top-K, parameters to be perturbed are carefully calculated. To investigate whether the random selection in Random-k is efficient enough or can be further improved, we designed a set of experiments in which $K$ in $k\%$ parameters selected by Random-k is replaced by those chosen by Top-K. This formed a mixed Random-TopK unlearning strategy, and related experiments are conducted using settings similar to the aforementioned ones. The $Acc_{RE}$ for Top-K, Random-k, and Random-TopK after 100 training epochs are shown in Fig.8.
Based on the $Acc_{RE}$ trends for Top-K, Random-k, and Random-TopK, at the beginning of training, the $Acc_{RE}$ for Top-K is higher than that for Random-k and Random-TopK. This implies that as the number of perturbed parameters in the model increases, more knowledge is eliminated. For Random-TopK, its $Acc_{RE}$ is lower than both Top-K and Random-k, indicating that over-modifying parameter based on those with the highest sensitivity will be more time-consuming. For models like ResNet18 (Fig.8(a)-8(d)), VGG (Fig.8(e)-8(h)), and DenseNet (Fig.8(i)-8(l)), Random-TopK still exhibits a certain upward trend in $Acc_{RE}$, suggesting that even with excessive parameter modifications, Random-TopK can still achieve partial unlearning effects. However, we notice that when using the Random-Topk stragegy on GoogLeNet (Fig.8(m)-8(p)), Random-TopK still exhibits a certain upward trend in $Acc_{RE}$, suggesting that even with excessive parameter modifications, Random-TopK can still achieve partial unlearning effects. However, we notice that when using the Random-Topk stragegy on GoogLeNet (Fig.8(m)-8(p)), the model fell into a state of complete non-learning. We believe this may be due to the different sensitivities and parameter dependencies between the different branches of the Inception module in GoogLeNet. Employing the Random-TopK strategy might introduce complexities, adversely affecting certain branches.
@@ IV-C5 Unlearning Degree Evaluation
Figure 9: Unlearning degree (%, ranges in [0, 90%]) evaluation for different unlearning strategies on ResNet18 using CIFAR-10 under different unlearning data ratio (5% - 20%). The higher unlearning degree is regarded as the better.
As aforementioned, in order to measure the unlearning degree for each unlearning strategy, we use Equation.13 to calculate the performance difference on the perturbed data $D_{p}$ generated by SPD-GAN between the source model $M$ and unlearning model $M_{UL}$. Utilizing source model $M$ and unlearning model $M_{UL}$ obtained from applying different unlearning strategies on CIFAR-10 with ResNet18 as the discriminator, the noise generator $G(\cdot)$ is trained.
The resulting unlearning degree is illustrated in Fig.9. The uniform random output accuracy for CIFAR-10 is around 10%, therefore, when we conduct experiments with CIFAR-10, the range of unlearning degree is [0, 90%]. The degree of unlearning indicates, beyond the model generalization caused high $Acc_{UL}$, the extent to which the machine unlearning actually eliminates the influence of unlearning data on the model. From the experimental results in Fig.9, it can be observed that Top-K achieved the best degree of unlearning (peaking at 88.4% when 10% data was unlearned), followed closely by Random-k, which reached a suboptimal level of unlearning (peaking at 86% when 5% of the data was unlearned). However, due to EU-K and CF-K freezing most of parameters and modifying the last $K$ layers, their degree of unlearning is significantly lower than that of Top-K and Random-k. Moreover, at lower ratio of unlearning data, such as 5%, the degree of unlearning across different unlearning strategies if relatively higher. This implies that as more data needs to be forgotten, achieving better unlearning effects may become more difficult.
@@ V Conclusion
In this paper, we propose fine-grained model weights perturbation methods to achieve inexact machine unlearning. Our method ensures a higher degree of unlearning the information from dataset $D_{UL}$ while maintaining model indistinguishability and achieving largest acceleration effect. Furthermore, we design a novel machine unlearning degree quantification method using SPD-GAN to break the i.i.d attribute of the unlearning data. The results further demonstrate that our proposed Top-K based inexact unlearning strategy achieves the best unlearning effect and reveal that it is more difficult to unlearn more data.
@@ Acknowledgments
The work is supported by the National Natural Science Foundation of China (Grant Nos. 62225205, 92055213), the National Key Research and Development Program of China (2021ZD40303), the Science and Technology Program of Changsha (kh2301011), Shenzhen Basic Research Project (Natural Science Foundation) (JCYJ20210324140002006) and China Scholarship Council.

### v4

@@ IV-C2 Determination of $K$ value for Top-K
$|D_{UL}|/|D|$ Strategy $Acc_{UL}$ $Acc_{RE}$ FR MRR Similarity Unlearn Time
5% Top-K 80.01% 97.61% 19.97% 97.61% 88.37% 76.97s
@@
We aim to achieve the effect of inexact machine unlearning by minimal number of parameter perturbed, to investigate the optimal value of $K$ for Top-K unlearning strategy, we conduct experiments with different values of $K$ within the range of 0 to 50 when the unlearning data ratio varies from 5% to 20%. Following the above settings, we apply Top-K to four models over 100 epochs. A metric to determine the optimal value of $K$ is necessary. Since machine unlearning aims to eliminate the influence of unlearning data while trying to not hurt the performance of remaining data, we calculate the accuracy difference between $Acc_{RE}$ and $Acc_{UL}$. As illustrated in Fig.5, the $K$ value corresponding to the maximum accuracy difference is identified as the optimal. As shown in Tab.I, considering the optimal $K$ values under four unlearning data ratios, we set $K=45$ for ResNet18 and DenseNet, $K=30$ for VGG, and $K=50$ for GoogLeNet.
As shown in Fig.5, although we have chosen the $K$ corresponding to the maximum accuracy difference, in actual experiments, the selection of $K$ values fluctuating within the range of 0 to 50 does not bring significant variability to the Top-K unlearning. For example, in Fig.5(a), for ResNet18, setting K=25 versus 45 does not make a substantial difference. Additionally, measuring the difference in accuracy is a sufficient but not necessary condition for the effectiveness of machine unlearning. In other words, for a fixed model, dynamically adjusting the value of $K$ based on the unlearning data ratio is highly impractical. Therefore, additional efforts or computational costs to determine a very precise $K$ value are unnecessary; we can roughly set it to a value such as 45.
IV-C3 Effectiveness of Top-K and Random-k
Fig.1 further shows the accuracy after applying the Random-k on the CIFAR-10 for 50 epochs. The accuracy of $D_{RE}$ is significantly higher than that of $D_{UL}$. As shown in Fig.6, it can be observed that the application of the Top-K achieves accuracy close to 95% ($Acc_{RE}$) across the four models, indicating a strong memory retention capability. Meanwhile, the accuracy difference between unlearning model on the $D_{RE}$ and the $D_{UL}$ (both in Fig.6 and Fig.1) shows that both the Top-K and Random-k achieve the unlearning effect more or less.
(a) 5% on ResNet18
(b) 10% on ResNet18
(c) 15% on ResNet18
(d) 20% on ResNet18
(e) 5% on VGG
(f) 10% on VGG
(g) 15% on VGG
(h) 20% on VGG
(i) 5% on DenseNet
(j) 10% on DenseNet
(k) 15% on DenseNet
(l) 20% on DenseNet
(m) 5% on GoogLeNet
(n) 10% on GoogLeNet
(o) 15% on GoogLeNet
(p) 20% on GoogLeNet
@@ IV-C3 Effectiveness of Top-K and Random-k
Tab.II provides a more comprehensive overview of the effects of different unlearning strategies at various unlearning ratios ($|D_{UL}|/|D|$). Across different unlearning ratios, Top-K consistently exhibits the lowest $Acc_{UL}$ values (80.14%, 75.48%, and 75.46%), along with corresponding forgetting rates (19.84%, 24.52%, and 24.54%). This suggests that, relative to other strategies, Top-K achieves the most effective reduction of the impact of $D_{UL}$ on the model when the unlearning ratio belongs to 10%, 15%, 20%. However, for the 5% unlearning ratio under Random-k, its corresponding $Acc_{RE}$ and MRR of 90.07% noticeably lag behind the results of other strategies. We believe this disparity is due to the instability of the strategy that employs random parameter perturbation, leading to inconsistent unlearning effects. In the meanwhile, across both $Acc_{RE}$ and MRR, the performance of $M_{UL}$ from nearly all unlearning strategies is comparable to that of retraining models, achieving close to 100% accuracy. As for similarity, there isn’t a significant distinction among different strategies, but this does indicate that model indistinguishability is maintained across diverse strategies.
As the ratio of $D_{UL}$ increases, the model becomes more attuned to the characteristics of the $D_{UL}$, leading to a further reduction in unlearning accuracy ($Acc_{UL}$) under an ideal unlearning strategy. However, as the ratio of $D_{RE}$ decreases, the model’s discriminative ability for $D_{RE}$ ($Acc_{RE}$) may also decline. As shown in Tab.II, across varying unlearning ratios, Top-K and Random-k exhibit comparable $Acc_{RE}$ within their respective settings, while as the unlearning ratio increases, the unlearning accuracy $Acc_{UL}$ continues to decrease (from 80.01% to 75.46% for Top-K).
For EU-K, fine-tuning the last $K$ layers to eliminate the influence of $D_{UL}$ allows the model to leverage the unaffected features from the frozen earlier layers for better predictions. At higher unlearning ratios, this strategy results in a decrease in $Acc_{UL}$ (from 90.68% to 84.48% for EU-5), while the performance on $D_{RE}$ remains stable.
(a) 5%
(b) 10%
(c) 15%
(d) 20%
Figure 9: Unlearning degree (%, ranges in [0, 90%] for CIFAR-10) evaluation for different unlearning strategies on ResNet18 using CIFAR-10 under different unlearning data ratio ((a)5%, (b)10%, (c)15% and (d)20%). The higher unlearning degree is regarded as the better.
In contrast, for CF-K, the parameters of the last $K$ layers are reinitialized, granting the model enhanced flexibility to relearn and mitigate the influence of unlearning data. This leads to a continuous decline in $Acc_{UL}$ (from 86.52% to 82.86% for CF-5). However, reinitializing and retraining the last $K$ layers negatively impacts the model’s performance on the $D_{RE}$. As reflected in Tab.II, under the CF-K strategy, $Acc_{RE}$ also shows a declining trend as the unlearning ratio increases (from 99.16% to 93.14% for CF-5).
@@
To further illustrate the acceleration effect of unlearning strategies on different models compared to the retraining method, we also conduct machine unlearning tasks on VGG, DenseNet, and GoogLeNet. Fig.7 shows the acceleration ratio of each unlearning strategy compared to retraining under different models and unlearning ratios. The acceleration effect of Top-K and Random-k is better than other strategies. Taking Fig.7(a) as an example, when the unlearning ratio is 5%, Top-K can achieve an acceleration of more than 15$\times$, while Random-k achieves an acceleration close to 9$\times$. The acceleration from EU-K and CF-K methods is not significant. As the unlearning ratio increases, the acceleration of Top-K and Random-k slows down. This is because, as more data is to be forgotten, the degree of modification to the model deepens. In conclusion, we can ascertain that the Top-K achieves optimal unlearning performance with minimal parameter perturbed, maintains considerable memory retention capability, and achieves the fastest acceleration effect.
@@ IV-C4 Mixed Top-K and Random-k
In Random-k, parameters are chosen randomly to be perturbed, while in Top-K, parameters to be perturbed are carefully calculated. When the value of k is quite small, it is likely that the unlearning process only modifies redundant neuron weights, failing to achieve effective unlearning. To investigate whether the random selection in Random-k is efficient enough or can be further improved, we designed a set of experiments in which $K$ in $k\%$ parameters selected by Random-k is replaced by those chosen by Top-K. This formed a mixed Random-TopK unlearning strategy, and related experiments are conducted using settings similar to the aforementioned ones. The $Acc_{RE}$ for Top-K, Random-k, and Random-TopK after 100 training epochs are shown in Fig.8.
Based on the $Acc_{RE}$ trends for Top-K, Random-k, and Random-TopK, at the beginning of training, the $Acc_{RE}$ for Top-K is higher than that for Random-k and Random-TopK. This implies that as the number of perturbed parameters in the model increases, more knowledge is eliminated. For Random-TopK, its $Acc_{RE}$ is lower than both Top-K and Random-k, indicating that over-modifying parameter based on those with the highest sensitivity will be more time-consuming. For models like ResNet18 (Fig.8(a)-8(d)), VGG (Fig.8(e)-8(h)), and DenseNet (Fig.8(i)-8(l)), Random-TopK still exhibits a certain upward trend in $Acc_{RE}$, suggesting that even with excessive parameter modifications, Random-TopK can still achieve partial unlearning effects. However, we notice that when using the Random-Topk stragegy on GoogLeNet (Fig.8(m)-8(p)), the model fell into a state of complete non-learning. We believe this may be due to the different sensitivities and parameter dependencies between the different branches of the Inception module in GoogLeNet. Employing the Random-TopK strategy might introduce complexities, adversely affecting certain branches.
@@ IV-C5 Unlearning Degree Evaluation
As aforementioned, in order to measure the unlearning degree for each unlearning strategy, we use Eq.14 to calculate the performance difference on the perturbed data $D_{p}$ generated by SPD-GAN between the source model $M$ and unlearning model $M_{UL}$. Utilizing source model $M$ and unlearning model $M_{UL}$ obtained from applying different unlearning strategies on CIFAR-10 with ResNet18 as the discriminator, the noise generator $G(\cdot)$ is trained.
The resulting unlearning degree is illustrated in Fig.9. The uniform random output accuracy for CIFAR-10 is around 10%, therefore, when we conduct experiments with CIFAR-10, the range of unlearning degree is [0, 90%]. The degree of unlearning indicates, beyond the model generalization caused high $Acc_{UL}$, the extent to which the machine unlearning actually eliminates the influence of unlearning data on the model. From the experimental results in Fig.9, it can be observed that Top-K achieved the best degree of unlearning (peaking at 88.36% when 10% data was unlearned), followed closely by Random-k, which reached a suboptimal level of unlearning (peaking at 86.04% when 5% of the data was unlearned). However, due to EU-K and CF-K freezing most of parameters and modifying the last $K$ layers, their degree of unlearning is significantly lower than that of Top-K and Random-k. Moreover, at lower ratio of unlearning data, such as 5%, the degree of unlearning across different unlearning strategies if relatively higher. As unlearning ratio increases, the unlearning effect gradually diminishes. This implies that as more data needs to be forgotten, achieving better unlearning effects may become more difficult.
@@ V Conclusion
In this paper, we propose fine-grained model weights perturbation methods to achieve inexact machine unlearning. Our method ensures a higher degree of unlearning the information from dataset $D_{UL}$ while maintaining model indistinguishability and achieving largest acceleration effect. Furthermore, we design a novel machine unlearning degree quantification method using SPD-GAN to break the i.i.d property of the unlearning data. The results further demonstrate that our proposed Top-K based inexact unlearning strategy achieves the best unlearning effect and reveal that it is more difficult to unlearn more data.
@@ Acknowledgments
The work is supported by the National Key Research and Development Program of China (2023YFB3002200), the National Natural Science Foundation of China (Grant Nos. 62225205, 92055213), the Science and Technology Program of Changsha (kh2301011), Shenzhen Basic Research Project (Natural Science Foundation) (JCYJ20210324140002006) and China Scholarship Council. Zhiwei Zuo did part of the work while she was a visiting student at NTU Singapore.